Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form 
y=mx+b.


(6,7); 
x+8y=7

Answer by Alan3354(69443)

Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form 
y=mx+b.


(6,7)
x+8y=7
-------------
Find the slope, m, of the line.
Then use y = mx + b and the point to find b.
Then you're done.